Circular No. 8210

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ij IN UGC 3336 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova on unfiltered CCD images taken on Sept. 25.954 (when the new object appeared at mag 18.3) and 26.862 UT (at mag 17.8) with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol. SN 2003ij is located at α = 5^{h}55^{m}41^{s}.18, δ = +85^{o}54′21′′.7, which is ≈ 3′′.7 west and 31′′.4 south of the center of UGC 3336. The new object is not present on Boles’ images from Aug. 29 and Sept. 15 (limiting mag 19.5), and it is not present on Palomar Sky Survey red or blue plates from 1996.

SUPERNOVA 2003ii IN MCG +06-6-53 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, write that a spectrum (range 370–750 nm) of SN 2003ii (cf. IAUC 8208), obtained by M. Calkins on Sept. 26.37 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Narrow emission lines superposed on the spectrum of the supernova indicate a recession velocity of 13400 km/s for the host galaxy; using this value, the supernova expansion velocity derived from the minimum of the Hβ line is 8600 km/s.

V838 MONOCEROTIS L. A. Crause, University of Cape Town, communicates post-conjunc-

tion Cousins photometry of V838 Mon, obtained with the South African Astronomical Observatory 1.0-m telescope on Sept. 11.13 UT: V = 15.65, V − R = +3.22, V − I = +6.17. The star brightened by ∼ 0.5 mag in V during the past 4 months, and the light echo has expanded to a radius of ∼ 52′′.

NOVAE IN M31 T. A. Rector reports the discovery by the University of Alaska (An-

chorage) Nova Search Team of an apparent nova in M31 on Hα images obtained with the WIYN 0.9-m telescope at Kitt Peak. The new object is located at α = 0^{h}42^{m}49^{s}.61, δ = +41^{o}18′01′′.8 (equinox 2000.0), and Rector provides the following Hα magnitudes: Aug. 3.83 UT, [19.0:; 31.78, 16.1; Sept. 17.31, 17.4; 19.31, 17.4. Rector also provides Hα magnitudes for the third apparent nova in M31 tabulated on IAUC 8165: Aug. 3.83, 14.8; 31.78, 16.1; Sept. 17.31, 17.0; 19.31, 17.0.

2003 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green